OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (1, 3) and is perpendicular to the line y = –x + 2?

OpenStudy (anonymous):

C=4? but thats not one of the multiple choice

OpenStudy (anonymous):

@JayDS

OpenStudy (anonymous):

y = x – 4 y = x + 2 y = –x + 2 y = –x – 4

OpenStudy (anonymous):

these are the choices

OpenStudy (anonymous):

sorry, my bad, I thought I read parallel, if it's perpendicular.

OpenStudy (anonymous):

slope-intercept form is y=mx + c m=gradient/slope c=y-intercept we are given a point (1,3) and the hint that it is perpendicular to the line y=-x + 2 when they say perpendicular we need to follow this rule. m1 x m2 = -1 m1 is the gradient from the equation they have given us which is -1 m2 is the gradient of the unknown equation we want to find. -1 x m2 = -1 m2 = -1/-1 m2 = 1 now substitute (1,3) and m=-1 into the basic formula. 3=1(1) + c c= ..... now re-write your whole formula with the gradient and y-intercept(the c value you calculated), and done.